# JiraiBench: A Bilingual Benchmark for Evaluating Large Language Models’ Detection of Human Self-Destructive Behavior Content in Jirai Community

### v2

@@ Appendix A Data Example
[…]
Table 3: Examples of Self-Destructive Content Categories in Chinese and Japanese
Appendix B Annotation Guidelines
B.1 Annotation Objectives and Framework
@@
    Drug Overdose (OD): Content containing drug overdose/misuse-related self-harm or extremely dangerous behaviors.
@@
    Eating Disorders (ED): Content containing eating disorder tendencies, such as anorexia, binge eating, or purging behaviors.
@@
    Self-harm or Suicidal Ideation (SH): Content containing self-injury, suicidal thoughts, or suicide plans.
For each category (OD, ED, SH), annotation values were assigned as follows:
@@
Note: If a post received a value of 0 across all three categories (OD, ED, SH), it was still retained in the dataset (indicating the post did not manifest any self-destructive/self-harm related content).
B.2 Annotation Procedure
B.2.1 Complete Reading
@@
B.2.2 Identification and Classification
  • Determine whether the post relates to any of the following categories: OD, ED, SH.
  • If multiple content types are mentioned (e.g., “excessive weight loss + self-harm”), each relevant category must be independently assessed and annotated.
@@ B.2.2 Identification and Classification
B.2.3 Distinguishing Between Poster vs. Third Party
@@
  • Completely lacking such content: If no suggestions or comments about OD, ED, or SH appear, mark as “0”.
B.2.4 Concurrent Annotation
  • As this employs multi-label annotation, assign values 0/1/2 to each of the three categories (OD, ED, SH), resulting in three annotation values.
B.3 Detailed Annotation Guidelines and Examples
@@ B.3.1 Drug Overdose (OD)
    • Posts completely lacking mention of drugs, narcotics, overdose, etc.
    • Example: “I’m in a good mood today, went for a run.”
  • 1: Poster’s own OD tendency
    • Posts mentioning the author’s own drug overdose or misuse behaviors, intentions, or plans.
@@
  • 2: Third-party OD comments
    • Posts mentioning others’ drug overdose or social incidents, not the author’s actions.
@@ B.3.2 Eating Disorder (ED)
    • Posts without any eating disorder-related expressions.
@@
B.3.3 Self-Harm or Suicidal Ideation (SH)
@@
    • Completely lacking any hint of self-harm, suicidal thoughts, or behaviors.
@@
  • 1: Poster’s own self-harm/suicidal tendency
    • Clearly mentioning the author’s own self-harm or suicidal intentions, behaviors, plans, or strong emotions.
@@
  • 2: Third-party self-harm/suicide comments
    • Posts only discussing others’ self-harm/suicide situations, social news, seeking help, etc.
@@
B.4 Common Issues and Handling Principles
B.4.1 Ambiguous/Vague Information
@@
B.4.2 Jokes, Metaphors, or Rhetoric
@@
B.4.3 Multiple Labels
  • A post may simultaneously contain two or three self-harm tendencies. For example, if a post discusses both drug misuse and self-harm tendencies, mark “OD=1, SH=1”. If ED is not mentioned, it remains 0.
@@ B.4.3 Multiple Labels
B.4.4 All Three Categories Marked as 0
  • This means the post contains no text or implications related to OD, ED, or SH, in which case the post should be retained with labels (OD=0, ED=0, SH=0).
B.5 Annotation Format
@@
B.6 Inter-annotator agreement
Table 4 presents the inter-annotator agreement metrics for the content annotation tasks related to drug overdose (OD), eating disorder (ED), and self-harm (SH). Two reliability measures are reported: Cohen’s Kappa and Fleiss’ Kappa, which are widely used to assess the agreement between annotators while accounting for chance agreement.
Cohen’s Kappa measures pairwise agreement between two annotators or an annotator and an expert. Higher values indicate stronger agreement, with scores above 0.6 typically interpreted as substantial agreement and scores above 0.8 as almost perfect agreement. In contrast, Fleiss’ Kappa generalizes Cohen’s Kappa to more than two annotators, providing an overall measure of agreement across all annotators.
For the overdose task (OD), the pairwise Cohen’s Kappa scores range from 0.6434 (A2 vs. Expert) to 0.7491 (A1 vs. A2), with an average of 0.6917 and a Fleiss’ Kappa of 0.6867. These values indicate substantial agreement among annotators for this task. The eating disorder task (ED) demonstrates higher reliability, with pairwise Cohen’s Kappa scores ranging from 0.7089 to 0.8498, resulting in an average Kappa of 0.7820 and a Fleiss’ Kappa of 0.7844. Similarly, the self-harm task (SH) achieves strong agreement, with pairwise Kappa values between 0.7122 and 0.8551, an average of 0.7785, and a Fleiss’ Kappa of 0.7813.
Overall, the results suggest moderate to substantial inter-annotator agreement across all tasks, with the strongest reliability observed in the eating disorder task. This indicates that the annotation guidelines were consistently applied by the annotators, supported by their expertise and familiarity with the content. The use of both Cohen’s and Fleiss’ Kappa provides a comprehensive evaluation of agreement, capturing both pairwise and overall consistency.
@@ B.6 Inter-annotator agreement
Appendix C Prompt
Table 5 illustrates the prompt templates designed for detecting self-destructive behavior across three languages—Chinese, Japanese, and English—and three behavior categories: OD, ED, and SH. These prompts are specifically crafted to leverage zero-shot and two-shot learning capabilities of LLMs, bypassing the need for traditional supervised training with labeled datasets.
@@ Appendix C Prompt
C.1 Zero-Shot and Two-Shot Use
@@
    Zero-shot settings: The prompts provide all necessary context and instructions within a single input, enabling the LLM to perform the classification task without requiring any prior task-specific examples. For example, in the overdose (OD) category, the prompt instructs the LLM to determine whether the content relates to drug abuse and to return a single numerical label as the result.
@@
C.2 Multilingual and Multi-Category Design
The templates are standardized across Chinese, Japanese, and English to ensure consistency in task execution regardless of the language. This multilingual alignment allows the LLM to operate effectively across diverse linguistic contexts without retraining or fine-tuning. Similarly, the three categories (OD, ED, and SH) are designed with uniform label definitions to ensure interpretability and comparability across behaviors.
@@ C.2 Multilingual and Multi-Category Design
Determine if the following social media content is related to drug abuse. Please output only one number as the result:
@@
1: Implies or directly mentions drug abuse
2: Description or comment on someone else’s drug abuse behavior
[…]
@@
Determine if the following social media content is related to self-harm or suicidal tendencies. Please output only one number as the result:
@@
1: Implies or directly mentions self-harm/suicide
2: Description or comment on someone else’s self-harm/suicide behavior
Table 5: Prompt templates used for self-destructive behavior detection across three languages (Chinese, Japanese, and English) and three categories: overdose (OD), eating disorders (ED), and self-harm (SH).
Appendix D JiraiLLM-Qwen

### v3

@@ Appendix B Data Example
Drug Misuse (DM) gbsn喜欢ymsf的濒死感。每次o都想写遗书给朋友留点谷。(I like the near-death feeling from taking Dextromethorphan. Every time I overdose, I want to write a suicide note and leave some final words to my friends.) minODもアムカも、やってみようかなくらいの気持ちだわ。 (I feel like I might try an overdose on amoxicillin (Amca), just considering it a bit)
Eating Disorder (ED) gbsn暴食后的催吐让人恶心,父母的控制欲令我窒息。 (Purging after binge eating makes me sick, and my over-controlling parents suffocate me.) min昨日から食べても食べても吐いてしまうくぽ… (Since yesterday, no matter how much I eat, I keep throwing up…")
Non-sucidal self injury (NSSI) gbsn每次划开手腕的时候,那种痛感反而让我感到平静。 (Every time I cut my wrist, the pain actually makes me feel calm.) min誰も気づかないように、太ももを切った。痛みだけが本当の感情。 (I cut my thigh where no one would notice. The pain is the only real feeling.)
Table 5: Paraphrased example of Risky health behavior Content Categories in Chinese and Japanese
Appendix C Annotation Guidelines
C.1 Annotation Objectives and Framework
@@
    Drug Misuse (DM): Content containing risky health behavior related to drug misuse or extremely dangerous behaviors.
@@
    Eating Disorders (ED): Content that contains eating disorder tendencies, such as anorexia, binge eating, or purging behaviors.
@@
    Non-Suicidal Self-Injury and Suicidal Ideation (NSSI/SI): Content containing nonsuicidal self-injury, suicidal thoughts, or suicide plans.
For each category (DM, ED, NSSI/SI), annotation values were assigned as follows:
@@
Note: If a post received a value of 0 across all three categories (DM, ED, NSSI/SI), it was still retained in the dataset (indicating the post did not manifest any risky health behavior-related content).
C.2 Annotation Procedure
C.2.1 Complete Reading
@@
C.2.2 Identification and Classification
  • Determine whether the post relates to any of the following categories: DM, ED, NSSI/SI.
  • If multiple content types are mentioned (e.g., “excessive weight loss + non-suicidal self-injury”), each relevant category must be independently assessed and annotated.
@@ C.2.2 Identification and Classification
C.2.3 Distinguishing Between Poster vs. Third Party
@@
  • Completely lacking such content: If no suggestions or comments about DM, ED, or NSSI/SI appear, mark as “0”.
C.2.4 Concurrent Annotation
  • As this employs multi-label annotation, assign values 0/1/2 to each of the three categories (DM, ED, NSSI/SI), resulting in three annotation values.
C.3 Detailed Annotation Guidelines and Examples
@@ C.3.1 Drug Misuse (DM)
    • Posts completely lacking mention of drugs, narcotics, misuse, etc.
    • Example: “I am in a good mood today, went for a run.”
  • 1: Poster’s own DM tendency
    • Posts mentioning the author’s own drug misuse behaviors, intentions, or plans.
@@
  • 2: DM comments from third parties
    • Posts mentioning others’ drug misuse or social incidents, not the author’s actions.
@@ C.3.2 Eating Disorders (ED)
    • Posts without any expressions related to eating disorders.
@@
C.3.3 Non-Suicidal Self-Injury and Suicidal Ideation (NSSI/SI)
@@
    • Completely lacking any hint of non-suicidal self-injury, suicidal thoughts, or behaviors.
@@
  • 1: Poster’s own NSSI/SI tendency
    • Clearly mentioning the author’s own non-suicidal self-injury or suicidal intentions, behaviors, plans, or strong emotions.
@@
  • 2: Third-party NSSI/SI comments
    • Posts only discussing others’ self-injury/suicide situations, social news, seeking help, etc.
@@
C.4 Common Issues and Handling Principles
C.4.1 Ambiguous/Vague Information
@@
C.4.2 Jokes, Metaphors, or Rhetoric
@@
C.4.3 Multiple Labels
  • A post may simultaneously contain two or three risky health behavior tendencies. For example, if a post discusses both drug misuse and self-injury tendencies, mark “DM=1, NSSI/SI=1”. If ED is not mentioned, it remains 0.
@@ C.4.3 Multiple Labels
C.4.4 All Three Categories Marked as 0
  • This means the post contains no text or implications related to DM, ED, or NSSI/SI, in which case the post should be retained with labels (DM=0, ED=0, NSSI/SI=0).
C.5 Annotation Format
@@
Appendix D Inter-annotator Agreement
@@
Appendix E Prompt
Table 7 illustrates the template templates designed to detect risky health behavior in three languages: Chinese, Japanese, and English, and three categories of behavior: DM, ED, and NSSI. These prompts are specifically designed to leverage the zero-shot and two-shot learning capabilities of LLMs, bypassing the need for traditional supervised training with labeled datasets.
@@ Appendix E Prompt
E.1 Zero-Shot and Two-Shot Use
@@
    Zero-shot settings: The prompts provide all necessary context and instructions within a single input, enabling the LLM to perform the classification task without requiring any prior task-specific examples. For example, in the drug misuse (DM) category, the prompt instructs the LLM to determine whether the content relates to drug abuse and to return a single numerical label as the result.
@@
E.2 Multilingual and Multi-Category Design
Templates are standardized in Chinese, Japanese, and English to ensure consistency in task execution regardless of the language. This multilingual alignment allows the LLM to operate effectively across diverse linguistic contexts without retraining or fine-tuning. Similarly, the three categories (DM, ED, and NSSI) are designed with uniform label definitions to ensure interpretability and comparability across behaviors.
@@ E.2 Multilingual and Multi-Category Design
Determine if the following social media content is related to drug misuse. Please output only one number as the result:
@@
1: Implies or directly mentions drug misuse
2: Description or comment on someone else’s drug misuse behavior
ED gbsn请判断以下社交媒体内容是否涉及进食障碍。请只输出一个数字作为结果:
@@
Determine if the following social media content is related to non-suicidal self-injury or suicidal tendencies. Please output only one number as the result:
@@
1: Implies or directly mentions non-suicidal self-injury/suicide
2: Description or comment on someone else’s non-suicidal self-injury/suicide behavior
Table 7: Prompt templates used for risky health behavior detection across three languages (Chinese, Japanese, and English) and three categories: drug misuse (DM), eating disorders (ED), and non-suicidal self-injury and suicidal ideation (NSSI/SI).
Appendix F JiraiLLM-Qwen